# Shuffled Linear Regression via Spectral Matching

### v1

@@ I Introduction
where ${\bf y}_{i}\in\mathbb{R}^{m}$ and ${\bf x}_{i}^{\star}\in\mathbb{R}^{n}$ are the $i$-th measurement vector and latent feature vector, respectively, ${\bf A}\in\mathbb{R}^{m\times n}$ is a known sensing matrix, $\boldsymbol{\Pi}^{\star}\in\mathcal{P}_{m}$ is an unknown permutation matrix from the set of $m\times m$ permutation matrices $\mathcal{P}_{m}$, and ${\bf n}_{i}$ is an additive measurement noise vector independent of ${\bf x}_{i}$.
@@ I-B Our Contributions
Motivated by these considerations, in this work, we explore spectral matching to solve the SLR problems under the large-system limit. We consider the setup where the number of measurements $t$ significantly exceeds the dimension of measurements $m$. Our approach utilizes measurement samples to estimate the eigenbases of the signal sample covariance matrix. We then determine the optimal permutation by aligning the spectral components through spectral matching. This alignment allows us to reformulate the problem into traditional least squares (LS) or LASSO optimization frameworks. Finally, we extend our method to address the practical challenge of three-dimensional (3D) image registration. The key contributions of our work are summarized as follows:
@@
    We propose spectral matching methods for shuffled LS and shuffled LASSO problems and demonstrating a reduced computational complexity compared to existing algorithms.
@@ II Problem Formulation
$(\widehat{\boldsymbol{\Pi}},\widehat{\bf X})\in\operatorname*{\mathop{\arg\min}}_{{\bf X}\in\mathbb{R}^{d\times t},\boldsymbol{\Pi}\in\mathcal{P}_{m}}\lVert\boldsymbol{\Pi}^{T}{\bf Y}-{\bf A}{\bf X}\rVert_{F}^{2}.$ (2)
@@ III Spectral Matching
In summary, our approach first computes the sample covariance matrix and its eigendecomposition by (7) and (8). Subsequently, we estimate $\widehat{\boldsymbol{\Pi}}$ by solving the linear assignment problem in (9). The solution for ${\bf X}$ is then derived either directly from (4) or by solving the resultant convex problem in (5). The total computational complexity for computing (7)–(9) scales at the rate of $\mathcal{O}(m^{2}(m+t))$. In contrast, the state-of-the-art approximate branch-and-bound algorithm in [SLR_ref6], applicable only when $t=1$, exhibits a prohibitive complexity of $\mathcal{O}(n^{2m})$ for large values of $n$. Additionally, the complexity of the spatial-domain matching algorithm in [pmlr-v119-zhang20n] grows at a rate of $\mathcal{O}(m^{2}(m+n+t))$.
@@ IV-A Permutation Estimation Error Analysis
We begin by examining the permutation estimation error for the solution in (9). Denote $k\triangleq|\mathcal{A}|\leq\min\{m,n\}$. To assess the accuracy of the estimate $\widehat{\boldsymbol{\Pi}}$, we introduce a noiseless permutation estimation problem as a counterpart to (9):

### v2

@@ I Introduction
where ${\bf y}_{i}\in\mathbb{R}^{m}$ and ${\bf x}_{i}^{\star}\in\mathbb{R}^{n}$ are the $i$-th measurement vector and latent feature vector, respectively, ${\bf A}\in\mathbb{R}^{m\times n}$ is a known sensing matrix, $\boldsymbol{\Pi}^{\star}\in\mathcal{P}_{m}$ is an unknown permutation matrix from the set of $m\times m$ permutation matrices $\mathcal{P}_{m}$, and ${\bf n}_{i}$ is an additive measurement noise vector independent of ${\bf x}_{i}^{\star}$.
@@ I-B Our Contributions
Motivated by these considerations, in this work, we explore spectral matching to solve the SLR problems under the large-system limit. We consider the setup where the number of measurements $t$ significantly exceeds the dimension of measurements $m$, i.e., $t\gg m$. Our approach utilizes measurement samples to estimate the eigenbases of the signal sample covariance matrix. We then determine the optimal permutation by aligning the spectral components through spectral matching. This alignment allows us to reformulate the problem into traditional least squares (LS) or LASSO optimization frameworks. Finally, we extend our method to address the practical challenge of three-dimensional (3D) image registration. The key contributions of our work are summarized as follows:
@@
    We propose spectral matching methods for shuffled LS and shuffled LASSO problems and demonstrate a reduced computational complexity compared to existing algorithms.
@@ II Problem Formulation
$(\widehat{\boldsymbol{\Pi}},\widehat{\bf X})\in\operatorname*{\mathop{\arg\min}}_{{\bf X}\in\mathbb{R}^{n\times t},\boldsymbol{\Pi}\in\mathcal{P}_{m}}\lVert\boldsymbol{\Pi}^{T}{\bf Y}-{\bf A}{\bf X}\rVert_{F}^{2}.$ (2)
@@ III Spectral Matching
Figure 1: Schematic view of the spectral matching method.
In summary, our approach first computes the sample covariance matrix and its eigendecomposition by (7) and (8). Subsequently, we estimate $\widehat{\boldsymbol{\Pi}}$ by solving the linear assignment problem in (9). The solution for ${\bf X}$ is then derived either directly from (4) or by solving the resultant convex problem in (5). The whole spectral matching process is summarized in Fig. 1.
The total computational complexity for computing (7)–(9) scales at the rate of $\mathcal{O}(m^{2}(m+t))$. In contrast, the state-of-the-art approximate branch-and-bound algorithm in [SLR_ref6], applicable only when $t=1$, exhibits a prohibitive complexity of $\mathcal{O}(n^{2m})$ for large values of $n$. Additionally, the complexity of the spatial-domain matching algorithm in [pmlr-v119-zhang20n] grows at a rate of $\mathcal{O}(m^{2}(m+n+t))$.
Remark 2.
We note that a strictly positive spectral gap $\delta$ in the selected eigencomponents is crucial for the effectiveness of our proposed approach. To achieve this, we prioritize the selection of principal eigencomponents that correspond to distinct eigenvalues in (10). This means that, in principle, our method works only if the matrix ${\bf A}{\bf C}_{X}{\bf A}$ has at least one eigencomponent corresponding to distinct and positive eigenvalues. In other words, the number of selected eigenvectors should be at least one, i.e., $\mathcal{A}\neq\emptyset$. We do acknowledge an extreme case that would violate this assumption: if all the eigenvalues of ${\bf A}{\bf C}_{X}{\bf A}$ are identical, i.e., when ${\bf A}{\bf C}_{X}{\bf A}$ is proportional to an identity matrix ${\bf I}_{m}$. In this case, it is impossible to distinguish the order from the spectrum.
@@ IV-A Permutation Estimation Error Analysis
We begin by examining the permutation estimation error for the solution in (9). We denote the number of selected eigenvectors in the considered linear assignment problem in (9) as
$k\triangleq|\mathcal{A}|\leq\min\{m,n\}.$ (12)
To assess the accuracy of the estimate $\widehat{\boldsymbol{\Pi}}$, we introduce a noiseless permutation estimation problem as a counterpart to (9):